OpenStudy (anonymous):

tan^2 theta = express in terms of sin theta only

OpenStudy (kc_kennylau):

This identity may be helpful: \(\cos^2\theta+\sin^2\theta=1\)

OpenStudy (***[isuru]***):

first can u write tan^2 theta using sin and cos ?

OpenStudy (***[isuru]***):

\[\tan^{2} \theta = \frac{ \sin^{2} \theta }{ \cos^{2} \theta}\]

OpenStudy (***[isuru]***):

now substitute the expression u can for cos^2 theta from the equation @kc_kennylau gave u ... can u do that ?

OpenStudy (kc_kennylau):

Looks like he won't respond...

OpenStudy (***[isuru]***):

yep :i

OpenStudy (kc_kennylau):

I'll come back when he responds :)